A wheel that is drawn by a belt is making 1revolution per second (r/s). If the wheel is 18cm in diameter, what is the linear velocity of the belt in cm/s? 


1
Expert's answer
2021-02-04T06:24:05-0500

diameter = 18 cm

radius = 9 cm

angular velocity ("\omega" ) = 1 rev/sec = (2"\pi") rad/sec

we know that, linear velocity (v) = angular velocity ("\omega") * radius (r)


v = "\omega" * r

v = 2"\pi" (rad/sec) * 9 (cm)

v = 18"\pi" cm/sec

v = 56.52 cm/sec